# The myth of declining competitive balance in the UEFA Champions League group stage

### v2

@@ 3.3 An illustrative example
The final group ranking is Real Madrid, Napoli, Braga, Union Berlin. Compared to the pot allocation, one discordant pair (Napoli and Real Madrid) exists, hence . This discordant pair affects the first two places in the group ranking, hence . Compared to the ranking implied by Elo ratings, there is one discordant pair (Braga and Union Berlin), hence . This discordant pair affects not only the first two places in the group ranking, hence .
@@ 3.4 Data
We consider all seasons between 2003/04 and 2023/24 when the basic structure of the UEFA Champions League has not changed. The pot allocation and group rankings are readily available from several sources; we have used Wikipedia after cross-checking with the official UEFA site.
Football Club Elo Ratings on a given date can be downloaded from api.clubelo.com/YYYY-MM-DD. In contrast to UEFA club coefficients, the Elo rating of a team changes during the season. Thus, the Elo ratings on 1 September are used for each season because the group stage is usually played between September and December. As usual, the Elo ratings are rounded to the nearest integer. Then Group B in the 2020/21 UEFA Champions League contains two teams (Shakhtar Donetsk and Borussia Mönchengladbach) with the same value (1766), which should be broken for indices and . The tie is decided by the higher value before rounding.
4 Results and discussion
@@
Figure 1 shows the evolution and the linear trend of all competitive balance measures. The linear trend has been obtained by estimating the following regression model:
where is the th competitive balance measure and is the error term in season (every season is denoted by its first year when the group stage is played). Furthermore, is the intercept and is the coefficient of the season.
The values of are always higher than , that is, competitive balance seems to be worse if all teams playing in the group are considered rather than only the three strongest teams. Consequently, the difference between the strengths of the weakest team and one of the three strongest teams is usually higher than the difference between the strengths of two of the three strongest teams. Regarding ex post competitive balance indices, their means are around 0.5 and even the smallest seasonal average is above 0.2. Therefore, both pot allocation and Elo ranking are better predictors of the final group ranking than a random permutation as expected. Group G in the 2019/20 UEFA Champions League is an interesting outlier, where the final ranking is the opposite of the ranking derived from the pot allocation.
@@
Figure 1 may suggest a small drop in competitive balance over the recent two decades since the slope of the linear trend is always positive. However, none of them are significant according to Table 2. The -values are much higher than the ones found by Triguero-Ruiz and Avila-Cano, (2023), which is below 0.02 for ex ante and approximately 0 for ex post competitive balance. Unsurprisingly, the values of are never above 0.1, in stark contrast to the results of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023).
@@
Nonetheless, Table 3 presents that some trends can be found in ex ante competitive balance. In particular, it declined until the 2014/15 season and improved between the 2014/15 and 2023/24 seasons. This pattern is less robust but still exists if only the three strongest teams are considered in all groups.
@@
How can these divergent conclusions be explained? Although both Ramchandani et al., (2023) (28 seasons from 1992/93 to 2019/20) and Triguero-Ruiz and Avila-Cano, (2023) (19 seasons from 1999/2000 to 2017/2018) study different samples, this cannot be the main reason. According to Ramchandani et al., (2023), the reduction in ex post competitive balance is even more evident between the 2002/03 and 2019/20 seasons, where they report for the linear regression (unfortunately, the -value is missing). Analogously, the five-year averages of the DCB index continuously increase (the competitive balance worsens) in both the ex ante and ex post settings (Triguero-Ruiz and Avila-Cano,, 2023).
@@
Our ex ante competitive balance measure is based on the more accurate Elo rating. Triguero-Ruiz and Avila-Cano, (2023) use the UEFA club coefficients, which have likely become more concentrated among the leading clubs without a parallel dominance in the Elo ratings. A potential reason might be that the best clubs in the top leagues—except England, where the set of competitors is wider—can somewhat relax in their domestic leagues now since the number of guaranteed slots in the UEFA Champions League group stage has increased from two to four for the four highest-ranked leagues. Therefore, they are able to increasingly focus on international competitions.
@@
Regarding ex post competitive balance, we compare the final group ranking to a benchmark provided by either the pot allocation or the Elo ranking. Since the HHI does not require such a benchmark, a worsening ex post competitive balance may be covered if our benchmark becomes a more accurate predictor over time. However, this is unlikely. Even though the seeding system has been reformed as discussed in Section 3, it has rather increased uncertainty (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). Similarly, the differences in the Elo ratings have not increased (Figure 1), thus, it would be difficult to assume that the quality of Elo ranking has changed. On the other hand, the clubs may focus in the group stage more strongly on their number of points than previously, possibly due to the increasing financial incentives provided by the UEFA. This leads to a higher HHI but a lower probability of losing stakeless games. Then the earlier studies have only mixed a favourable trend (teams do not lose their stakeless matches) with declining ex post competitive balance. The hypothesis may be tested in the future by an empirical investigation of stakeless games, which can be identified by the approach of Csató et al., (2024).
@@ 5 Conclusions
UEFA has chosen a risky strategy by fundamentally reforming the well-established group stage of its major tournament, the UEFA Champions League, from the 2024/25 season. According to the existing literature using convenient measures, UEFA should have acted effectively because of the notable drop in both the ex ante and ex post competitive balance of the group stage over the last two decades. Our paper has proposed six alternative but reasonable indices to check the robustness of these results. Crucially, no evidence is found for a declining competitive balance between the 2003/04 and 2023/24 seasons. Therefore, we urge researchers and analysts to consider a broader set of competitive balance measures in the future before proposing powerful policy interventions.

### v4

@@ 4.3 An illustrative example
Consequently, the stronger team wins a group match with a probability of 69% in average.
@@
Consequently, the stronger team wins a group match with a probability of 64% in average if we focus on the three favourites.
@@
The final group ranking is Real Madrid, Napoli, Braga, Union Berlin. Compared to the pot allocation, one discordant pair (Napoli and Real Madrid) exists, hence . However, this discordant pair affects the first two places in the group ranking, hence . Compared to the ranking implied by Elo ratings, there is one discordant pair (Braga and Union Berlin), hence . This discordant pair affects not only the first two places in the group ranking, hence .
@@ 4.4 Data
We consider all seasons between 2003/04 and 2023/24 when the tournament format of the UEFA Champions League did not changed. The pot allocation and group rankings are readily available from several sources; we have used Wikipedia after cross-checking with the official UEFA site.
Football Club Elo Ratings on a given date can be downloaded from api.clubelo.com/YYYY-MM-DD. In contrast to UEFA club coefficients, the Elo rating of a team changes during the season. Thus, the Elo ratings on 1 September are used for each season because the group stage is usually played between September and December. The Elo ratings are rounded to the nearest integer. Then Group B in the 2020/21 UEFA Champions League contains two teams (Shakhtar Donetsk and Borussia Mönchengladbach) with the same value (1766), which should be broken for indices and . It is decided by the higher value before rounding, but the tie-breaking does not affect our qualitative results.
4.5 The evolution of competitive balance
The approach of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023) is followed directly. For all the six (two ex ante and four ex post) indicators of competitive balance, the linear trend is obtained by estimating the following regression model:
where is the th competitive balance measure and is the error term in season (). Every season is denoted by its first year when the group stage is played. Furthermore, is the intercept and is the coefficient of the season. Competitive balance declined/did not changed/improved if is significantly positive/statistically zero/significantly negative.
4.6 Overview
Our results can be replicated by the following procedure:
1.
Collecting the publicly available Football Club Elo Ratings, UEFA Champions League pot allocations and group rankings between the 2003/04 and 2023/24 seasons (see Section 4.4 for details);
- 2.
3.
Estimation of linear regressions to check the existence of any trend in the seasonal averages of , , and – (see Section 4.5 for details).
5 Results
@@
Figure 1 shows the evolution and the linear trend of the six competitive balance indices. The means of ex post competitive balance measures are around 0.5, and even the smallest seasonal average is above 0.2. Therefore, both pot allocation and Elo ranking are better predictors of the final group ranking than a random permutation (Groot,, 2008) as expected. Group G in the 2019/20 UEFA Champions League is an interesting outlier, where the final ranking is the opposite of the ranking implied by the pot allocation.
@@
Figure 1 may suggest a small drop in competitive balance over the recent decades since the slope of the linear trend is always positive. However, none of them are significant according to Table 2. This is in stark contrast with Triguero-Ruiz and Avila-Cano, (2023), where the -values are below 0.02 for ex ante and approximately 0 for ex post competitive balance, respectively. Unsurprisingly, never exceeds 0.1, contrary to the results of Ramchandani et al., (2023) and Triguero-Ruiz and Avila-Cano, (2023).
@@
Nonetheless, Table 3 presents that some trends can be found in ex ante competitive balance. In particular, it declined until the 2014/15 season and improved between the 2014/15 and 2023/24 seasons. The pattern is less robust but still exists if only the three strongest teams are considered in all groups as index does.
6 Discussion
@@
How can these divergent conclusions be explained? Although both Ramchandani et al., (2023) (28 seasons from 1992/93 to 2019/20) and Triguero-Ruiz and Avila-Cano, (2023) (19 seasons from 1999/2000 to 2017/2018) study different samples, this cannot be the main reason. According to Ramchandani et al., (2023), the reduction in ex post competitive balance is even more evident between the 2002/03 and 2019/20 seasons, where they report for the linear regression; unfortunately, the -value is not reported. Analogously, the five-year averages of the DCB index continuously increase—the competitive balance worsens—in both the ex ante and ex post settings (Triguero-Ruiz and Avila-Cano,, 2023). Nonetheless, Table 3 suggests that the improvement in the recent seasons not considered in previous studies is partially responsible for the insignificant trend found here.
@@
Our ex ante competitive balance indicators are based on the more accurate Elo rating. Triguero-Ruiz and Avila-Cano, (2023) use the UEFA club coefficients, which have likely become more concentrated among the leading clubs without a parallel dominance in the Elo ratings. A potential reason might be that the best clubs in the top leagues—except England, where the set of competitors is wider—can somewhat relax in their domestic leagues now since the number of guaranteed slots in the UEFA Champions League group stage has increased from two to four for the four highest-ranked leagues. Therefore, they are able to increasingly focus on international competitions.
@@
Regarding ex post competitive balance, we compare the final group ranking to a benchmark provided by either the pot allocation or the Elo ranking. Since the does not require such a benchmark, a worsening ex post competitive balance may be covered if our benchmark becomes a more accurate predictor over time. However, this is improbable. Even though the seeding system has been reformed as discussed in Section 3, it has rather increased uncertainty (Corona et al.,, 2019; Dagaev and Rudyak,, 2019). Similarly, the differences in the Elo ratings have not increased (Figure 1), thus, the quality of Elo ranking seems to be unchanged. On the other hand, the clubs may focus in the group stage more strongly on their number of points than previously, possibly due to the increasing financial incentives provided by the UEFA. This leads to a higher and a lower probability of losing stakeless games. Then the earlier studies have only mixed a favourable trend (teams do not lose their stakeless matches) with declining ex post competitive balance. The hypothesis may be tested by an empirical investigation of stakeless games, which can be identified, for example, by the approach of Csató et al., (2024). Finally, it should be noted that the sample size is rather small. However, the novel format of the UEFA Champions League from the 2024/25 season does not allow for more observations.
@@ 7 Conclusions
UEFA has chosen a risky strategy by fundamentally reforming the well-established group stage of its major tournament, the UEFA Champions League, from the 2024/25 season. According to the existing literature based on straightforward -based measures, UEFA should have acted effectively because of the notable drop in both the ex ante and ex post competitive balance of the group stage over the last two decades. Our paper has proposed six alternative but reasonable indices to see the robustness of these results. Crucially, no evidence is found for a declining competitive balance between the 2003/04 and 2023/24 seasons. Therefore, we urge researchers and analysts to consider a broader set of competitive balance indicators in the future before proposing powerful policy interventions.
@@ Acknowledgements
One anonymous reviewer provided useful comments on an earlier draft.